OpenStudy (anonymous):

g(a)=3a+1 h(a)=a^2 -a find ( g/h) (a)

OpenStudy (anonymous):

I came out with -3a^2 + 1

hero (hero):

\[\frac{g}{h}(a) = \frac{g(a)}{h(a)}=\frac{3a + 1}{a^2 - a} = \frac{3a + 1}{a(a - 1)}\] I doubt if you can simplify any more than that

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):